OpenStudy (anonymous):

find the derivative f(x)=ln(3x-2)^8

jimthompson5910 (jim_thompson5910):

is that [ln(3x-2)]^8 or is it ln[ (3x-2)^8 ]

OpenStudy (anonymous):

\[f(x)=\ln(3x-2)^{8}\]

jimthompson5910 (jim_thompson5910):

hmm I'll go with the second one, but unfortunately what they gave you is a bit vague

jimthompson5910 (jim_thompson5910):

y = ln[ (3x-2)^8 ] can be written as y = 8*ln[ (3x-2) ] or just y = 8*ln(3x-2)

jimthompson5910 (jim_thompson5910):

so let's derive y = 8*ln(3x-2) we pull the constant out and ignore it then we derive ln(3x-2) to get 1/(3x-2) but remember you have to use the change rule, so you're multiplying that by the derivative of what's inside: which is 3x-2....so derive 3x-2 to get 3 and multiply it with 1/(3x-2) to get 3/(3x-2)

jimthompson5910 (jim_thompson5910):

so y = 8*ln(3x-2) turns into y ' = 8*(3/(3x-2)) then you just simplify to get y ' = 24/(3x-2)

OpenStudy (anonymous):

can you explain why you can pull the 8 out front please?

jimthompson5910 (jim_thompson5910):

because deriving something like y = 10*ln(x) or y = 20*ln(x) involves the same steps: pull out the constant, derive ln(x) to get 1/x, then reintroduce the constant back in

jimthompson5910 (jim_thompson5910):

it's based on the rule that if y = c*f(x), then y ' = c * f ' (x) where c is any constant

OpenStudy (anonymous):

Alright, this derivative stuff drives me crazy

OpenStudy (anonymous):

i understand that, but why can you pull the 8 from the exponent in front to make it a product?

jimthompson5910 (jim_thompson5910):

yeah it'll derive any one crazy...

jimthompson5910 (jim_thompson5910):

oh I'm using the rule that ln(x^y) = y*ln(x)

jimthompson5910 (jim_thompson5910):

that's how I went from y = ln[ (3x-2)^8 ] to y = 8*ln(3x-2)

OpenStudy (anonymous):

ahhh cool thanks!

jimthompson5910 (jim_thompson5910):

np

jimthompson5910 (jim_thompson5910):

thanks lol

OpenStudy (anonymous):

no thank you

jimthompson5910 (jim_thompson5910):

I'm glad it's making sense now

jimthompson5910 (jim_thompson5910):

I see, so you'll have to somehow connect the dots in your head so you can remember it all, but I'm sure you can definitely do that